### v1

# Effect of freestream turbulence on the coherent dynamics of a wind turbine wake

Abstract
The wake of a model wind turbine exposed to freestream turbulence (FST) with a variety of turbulent characteristics is studied through Particle Image Velocimetry experiments. The FST cases spanned turbulent intensities ($T_{i}$) in the range $1\%\lesssim T_{i}\lesssim 13.1\%$ and integral length scales ($L_{v}$) in the range $0.01D\lesssim L_{v}\lesssim 0.11D$ (where $D$ is the diameter of the turbine model). Increasing $T_{i}$ and $L_{v}$ in this range resulted in an earlier breakdown of the tip vortices which in turn resulted in an earlier onset of wake recovery. Typically, the amplitude of wake meandering increased in the presence of FST, but not monotonically with either $T_{i}$ or $L_{v}$. The nature of the low frequency region (typical of wake meandering frequencies) of the power spectra changed from a more discrete to a more broadband one as FST level was increased, highlighting a fundamental change in the nature of wake meandering with/without FST. Deeper insights into the origins, and subsequent evolution, of the various coherent motions (characterised by particular frequencies) in the presence of FST are obtained through analysis of the multi-scale triple-decomposed coherent kinetic energy budgets. The wake meandering modes in the presence of FST are shown to better utilize the mean velocity shear, extracting more energy from the mean flow while other sources such as non-linear triadic interactions and diffusion also gain in importance.
keywords:
Wind turbines, wakes, coherent structures
1 Introduction
Wind turbine wakes exhibit a large range of coherent structures such as the tip/root vortices, distinct vortex sheddings from the tower or the nacelle and wake meandering [Sherry et al., 2013, Okulov et al., 2014, De Cillis et al., 2021, Biswas and Buxton, 2024a]. These coherent structures play a critical role in determining the spatio-temporal evolution of the wake of a turbine [Lignarolo et al., 2015, Porté-Agel et al., 2020, Biswas and Buxton, 2024a, b] and hence, can have large impacts on farm-level performance [Vermeer et al., 2003, Barthelmie et al., 2007, Sanderse et al., 2011, Stevens and Meneveau, 2017]. The dynamics of the coherent structures depends on several factors including (a) the geometry of the turbine, such as the design of the blades [Abraham and Leweke, 2023, Dong et al., 2023], (b) the operating conditions, such as the tip speed ratio $\lambda$, (defined as $\lambda=\Omega R/U_{\infty}$, where $\Omega$ is the turbine’s rotational speed, $R$ is the turbine radius and $U_{\infty}$ is the freestream velocity) [Sherry et al., 2013, Okulov et al., 2014, Biswas and Buxton, 2024a, b] and (c) freestream conditions such as the freestream turbulence (FST) level [Mao and Sørensen, 2018, Hodgson et al., 2023, Bourhis et al., 2025].
Real wind turbines are almost always exposed to some level of FST. FST can be introduced for instance, due to the presence of the atmospheric boundary layer which is characterized by a broad range of scales due to the inherently large Reynolds number, or due to the presence of an upstream obstacle such as another wind turbine [Porté-Agel et al., 2020, Kosović et al., 2025]. The nature and the level of turbulence in an atmospheric boundary layer is linked to the boundary layer characteristics such as the boundary layer thickness and atmospheric stability [Peña et al., 2016, Vahidi and Porté-Agel, 2024, Kosović et al., 2025]. The level of FST is often measured using the turbulence intensity ($T_{i}$), defined as $T_{i}=\sqrt{(\overline{u^{\prime 2}})}/U_{\infty}$ (where $\sqrt{(\overline{u^{\prime 2}})}$ is the standard deviation of the velocity time history). In onshore sites wind turbines are exposed to higher $T_{i}$ ($T_{i}<15\%$) for most of their operating life [Elliott and Cadogan, 1990, Wagner et al., 2011, Peña et al., 2016], while for offshore sites freestream $T_{i}$ is generally milder: $6\%<T_{i}<8\%$ [Barthelmie et al., 2005, Türk and Emeis, 2010], due to the reduced surface roughness.
The effect of FST on the performance of a wind turbine is multifaceted. FST can impart fluctuating loads [Eggers Jr et al., 2003, Stanislawski et al., 2023, Robertson et al., 2019, de Oliveira et al., 2025] and significantly alter the power output from a turbine [Chamorro et al., 2013, Maganga et al., 2010, Blackmore et al., 2016]. Not only can it alter the performance of a standalone turbine, it can modify the spatio-temporal evolution of the wake, which can significantly change farm level performance [Porté-Agel et al., 2020]. On one hand, FST can make the wake meander more in the lateral direction [Espana et al., 2012, Porté-Agel et al., 2020], potentially reaching other turbines in a staggered arrangement, imparting harmful unsteady loads [Larsen et al., 2008]. On the other hand, FST can help re-energise the wake by enhancing lateral/vertical momentum transport from the freestream which can be beneficial for the downwind turbines in terms of power production [Gambuzza and Ganapathisubramani, 2023, Vahidi and Porté-Agel, 2024].
The effect of FST has been studied through a combination of numerical simulations (largely using large eddy simulation, LES), laboratory experiments using a small-scale representative wind turbine, and field experiments. High levels of FST has been shown to break down the tip vortices in the near wake [Chamorro and Porté-Agel, 2009, Wu and Porté-Agel, 2012]. These vortices act as a ‘shield’ in the near wake, inhibiting the exchange of mass, momentum and kinetic energy [Medici, 2005]. Lignarolo et al. [2015] performed particle image velocimetry (PIV) experiments on a two-bladed wind turbine model and showed that there was a net kinetic energy entrainment into the wake only after the tip vortices broke down. The recent study by Biswas and Buxton [2024b] reached a similar conclusion for a three-bladed wind turbine model. Therefore, high FST levels can accelerate wake recovery by breaking down this shield of vortices. In fact, a linear relationship between $T_{i}$ and the wake recovery rate has been reported [Niayifar and Porté-Agel, 2016, Carbajo Fuertes et al., 2018, Gambuzza and Ganapathisubramani, 2023].
In the near wake, freestream turbulence breaks down the dominant coherent structures i.e. the tip vortices. In contrast, freestream turbulence likely plays a different role on the dominant far wake structures, which have long been primarily associated with wake meandering. Although the dominance of wake meandering in the far wake has been known for many years, there still remain varied opinions on the genesis of the meandering motion. For instance, a distinct wake meandering frequency has been reported when freestream turbulence intensity was negligible [Okulov et al., 2014, Biswas and Buxton, 2024a] thereby associating wake meandering to a rotor scale bluff body shedding type instability [Medici and Alfredsson, 2008, Howard et al., 2015, Biswas and Buxton, 2024a]. In contrast, in other works wake meandering has been treated as a ‘random’ unsteady oscillation of the far wake driven by freestream turbulence [Porté-Agel et al., 2020]. Here, the wind turbine wake is thought to be passively advected by the large scale structures present in the freestream turbulence [Larsen et al., 2008]. Unifying the two schools of thoughts, there is now a growing evidence that in real wind turbines exposed to FST, the nature and intensity of wake meandering depends on multiple factors including the turbine geometry, operating condition and freestream conditions [Howard et al., 2015, Mao and Sørensen, 2018, Heisel et al., 2018, Gupta and Wan, 2019, Gambuzza and Ganapathisubramani, 2023, Biswas and Buxton, 2024a].
Although, the effect of $T_{i}$ on a turbine wake has been studied extensively, the need to characterize the spectral content of the inflow turbulence has been realized only recently [Mao and Sørensen, 2018, Gupta and Wan, 2019, Vinod and Banerjee, 2019, Gambuzza and Ganapathisubramani, 2023, Hodgson et al., 2023, Bourhis et al., 2025]. In engineering wake models, the growth of the wake is often assumed to be a function of turbulence intensity alone and the spectral characteristic of the incoming flow is often ignored [Gambuzza and Ganapathisubramani, 2023]. Low frequencies present in the inflow have been shown to modulate the instantaneous power output of a stand alone turbine [Chamorro et al., 2013]. There is a newly established consensus that wind turbines do not simply advect the structures present in the freestream as was originally assumed by the dynamic wake meandering model [Larsen et al., 2008]. Instead, a turbine has a dynamic response i.e. certain frequencies in the freestream can be amplified in the wake [Mao and Sørensen, 2018, Heisel et al., 2018, Gupta and Wan, 2019, Hodgson et al., 2023]. Mao and Sørensen [2018] identified nonlinear optimal perturbations to a wind turbine wake modeled as an actuator disk. They reported that as the inflow perturbation amplitude increased, the optimal perturbation frequency reduced from $St=0.64$ to $St=0.25$ for $Re_{D}=1000$. As $Re_{D}$ was increased to 3000, the optimal perturbation frequency further reduced to $St=0.16$. These low frequencies fall well within the accepted range of wake meandering frequencies [Medici and Alfredsson, 2008, Chamorro et al., 2013, Okulov et al., 2014]. Beyond $Re=3000$, Mao and Sørensen [2018] noted that the optimal frequency did not change, indicating a possible $Re$ invariance at all application specific Reynolds numbers. Similarly, through field experiments, Heisel et al. [2018] reported that the wind turbine amplifies a range of frequencies in the inflow yielding spectral peaks in the range $St=0.3-0.4$ in the far wake. Similarly, larger freestream characteristic length scales were shown to induce a stronger wake meandering and were linked to a faster kinetic energy entrainment from the freestream [Blackmore et al., 2014, Vahidi and Porté-Agel, 2024]. However, these studies modeled the turbine as an actuator disk so, the effect of the tip vortices was not considered. The results of these studies likely hold better in the far wake of a turbine which is better approximated by the wake of an actuator disk. Interestingly, Hodgson et al. [2023] reported a shorter near wake region and faster wake recovery for shorter integral time scales. These results indicate that the freestream scales may have a different effect in the near and the far wake [Hodgson et al., 2023]. The current study aims to further understand the effect of the freestream characteristic length scales on the near wake of a turbine.
Unlike several previous works where the statistical properties of the wake were studied in the presence of FST, we primarily focus on the dynamic properties. A series of time-resolved PIV experiments are performed on a wind turbine model exposed to freestream turbulence having different turbulence intensities ($T_{i}$) and integral length scales ($L_{v}$). In this work we build on our previous works by including freestream turbulence and we attempt to better understand the following aspects: (a) How does the energy content at the dominant frequencies in the near-wake change with freestream turbulence? (b) In the absence of freestream turbulence, Biswas and Buxton [2024a] found a scaling between the wake meandering frequency and $\lambda$ indicating that wake meandering could be a rotor scale bluff body shedding type instability. How does the nature of wake meandering change in the presence of freestream turbulence? (c) How does freestream turbulence affect the energy exchanges between the various coherent modes? (d) Finally, in all of the above what are the relative roles of $T_{i}$ and $L_{v}$?
Figure 1: (a) Parameter space of turbulence intensity ($T_{i}$) and integral length scale ($L_{v}$) produced by different turbulence generating grids. The error bars show the maximum and minimum values observed within the rotor diameter. (b) shows the effective tip speed ratios ($\lambda_{eff}$) for $\lambda_{\infty}=6$. $\lambda_{eff}$ is lower than $\lambda_{\infty}$ as the freestream velocity increases due to the blockage caused by the turbulence generating grids. The uncertainties in $\lambda_{eff}$ are due to the variation of the freestream velocity within the rotor diameter.
2 Methodology
Particle image velocimetry (PIV) experiments were conducted in the hydrodynamics flume in the Department of Aeronautics at Imperial College London. The flume had a cross-sectional area of $60\times 60$ cm2 at the operating water depth. The turbine radius ($R$) was 0.1 m and the model was the same as that detailed in [Biswas and Buxton, 2024a, b]. The model had a nacelle and tower associated with it such that it resembled a utility-scale turbine. The free-stream velocity ($U_{\infty}$) was kept constant at 0.2 ms-1. Experiments were performed at a single tip speed ratio $\lambda=6$. The Reynolds number based on the turbine diameter $Re_{D}$ (where $D$ is the turbine diameter) was $\approx 40000$.
Different freestream turbulence cases were obtained by using different passive turbulence generating grids similar to that used in our previous studies [Kankanwadi and Buxton, 2020, Chen and Buxton, 2023]. Separate PIV experiments were conducted to characterise the variation of turbulence intensity ($T_{i}$), integral length scales ($L_{v}$) and the mean streamwise velocity as a function of streamwise distance from the grids. The details of these experiments can be found in appendix 1. The parametric space (spanned by $T_{i}$ and $L_{v}$) in this study is shown in figure 1(a). Specific symbols are used for certain types of grids. The symbols represent the value of $T_{i}$ and $L_{v}$ at hub height, while the error bars show the maximum and minimum values observed within the rotor diameter (see the transverse variations of $T_{i}$ and $L_{v}$ in figures 23 and 25 respectively in appendix 1).
[…]
The grids also introduce a blockage to the freestream, resulting in an acceleration of the incoming flow. Although the tip speed ratio based on $U_{\infty}$ (denoted as $\lambda_{\infty}$) was kept constant at 6 for all the FST cases, due to the blockage induced increment of the freestream velocity, the effective tip speed ratio ($\lambda_{eff}$) reduced for all the FST cases. $\lambda_{eff}$ for the different cases is shown in figure 1(b). The symbols represent the values at the geometrical centreline whilst, the uncertainties are due to the transverse variation of the mean streamwise velocity. The uncertainties increase with $T_{i}$ as can be expected. The test cases are divided into 4 subgroups, primarily based on turbulence intensity. Group 1 has negligible $T_{i}$ while groups 2, 3 and 4 have low, mild and high $T_{i}$ respectively. This classification was done by comparing $T_{i}$ as well as the nature of the tip vortex related frequencies observed in the near wake as will be discussed in the upcoming sections. Nevertheless, hard boundaries likely do not exist between the adjacent subgroups.
Figure 2: Exemplar instantaneous vorticity fields for the cases (a) 1, (b) 2a, (c) 3, and (d) 4b.
The experiments with the wind turbine had the same field of view as that in experiment 1A discussed in [Biswas and Buxton, 2024b]. Therefore, the details of the experiments are only briefly discussed here. Only planar PIV experiments were performed at the diametrical plane of the rotor aligned with the tower’s axis. The experiments were time resolved. The acquisition frequency ($100$ Hz) was much higher than the turbine’s rotational frequency ($\approx 1.87$ Hz). The spatial resolution of the experiments ($\approx 0.0115D$) was good enough to capture the tip vortices (having size $\approx 0.07D$).
3 Vorticity field
Instantaneous vorticity fields for 4 cases, one from each subgroup (increasing turbulence intensity) are shown in figure 2 for $\lambda_{\infty}=6$ (see also supplementary videos 1-4). For case 1, the tip vortices undergo merging at $x/D\approx 1.5$ similarly to the observation in our earlier study (in the absence of FST) through a different set of experiments [Biswas and Buxton, 2024a]. Due to the outer fluid being non-turbulent, the wake edge is clearly marked by vorticity in case 1. The solid black line shows the mean wake centreline $y=y_{c}$ obtained by averaging over instantaneous wake centrelines (the loci of the minima in the instantaneous streamwise velocity profiles). The centreline bends towards the $y<0$ side due to the presence of the tower as was also discussed in previous studies [Pierella and Sætran, 2017, Biswas and Buxton, 2024a]. For low $T_{i}$ (case 2), tip vortices and their merging is still observed, however the trajectories of both the individual and merged vortices appear more erratic similar to the observation of Gambuzza and Ganapathisubramani [2023]. Large scale structures from the freestream penetrate deep into the wake, close to the wake centreline (see the supplementary video 2). As $T_{i}$ is increased further, the tip vortices break down much closer to the turbine as reported in previous studies [Chamorro and Porté-Agel, 2009, Wu and Porté-Agel, 2012, Gambuzza and Ganapathisubramani, 2023]. The presence of background turbulence makes it difficult to visually identify the wake edge for both medium and high $T_{i}$ cases.
Figure 3: Spectra at $x/D=0.5$ and $y/D=0.55$ for the cases (a) 1, (b) 2a, (c) 3, and (d) 4b.
4 Effect of freestream turbulence on the important frequencies
In order to understand the effect of FST on the coherent motions at different frequencies, the transverse velocity spectra (obtained using FFT) in the tip region of the near wake ($x/D=0.5$, $y/D=0.55$) are first shown in figure 3(a-d). The symbols in the top right corner correspond to the different FST cases shown in figure 1(a). For all the FST cases, the dominant frequency is $3f_{r}$, where $f_{r}$ is the turbine’s rotational frequency. $3f_{r}$ therefore denotes the frequency of the tip vortices for a three-bladed turbine. Apart from $3f_{r}$, $f_{r}$ is also found to be important and is observed in all cases. Biswas and Buxton [2024a] reported multiple harmonics of $f_{r}$ in the tip region in the absence of freestream turbulence. These harmonics up to $6f_{r}$ and even beyond are clearly observed for case 1. The orange band in the low frequency region $0.15\leq St_{D}\leq 0.4$ shows the typical frequency range of wake meandering reported in previous studies [Medici and Alfredsson, 2008, Okulov et al., 2014, Chamorro et al., 2013]. Note from figure 3(a) that without any freestream turbulence, the spectral energy content in this wake meandering frequency band (WMFB) is negligible in the tip region of the near wake. Contrastingly, for all the FST cases (case 2-4) non-negligible spectral peaks are observed in the WMFB even in the tip region. Especially for case 2a, there is a clear peak at $St_{D}\approx 0.27$. This difference indicates that the the physics underpinning wake meandering for case 1 are likely different from the cases with significant levels of freestream turbulence.
The role of the nacelle in the genesis of wake meandering has recently received some attention [Howard et al., 2015, Biswas and Buxton, 2024a]. Howard et al. [2015] argued that the nacelle’s vortex shedding provides an important perturbation that can trigger wake meandering in the absence of freestream turbulence. Biswas and Buxton [2024b] studied the nature of the wake meandering modes without FST. The wake meandering mode was shown to grow radially with downstream distance and it reached the tip region only after $x\approx 3D$. This justifies the absence of spectral peaks in the WMFB observed in figure 3(a). In the presence of freestream turbulence (cases 2-4), perturbations are available everywhere, unlike case 1 where the nacelle’s vortex shedding is the main source of the perturbation [Howard et al., 2015]. Therefore, with FST, wake meandering is likely to start closer to the turbine. As a result, a non-negligible amount of energy content is present in the WMFB for the FST cases. This will be discussed in more detail in the following paragraphs.
For low $T_{i}$ (case 2a), most of the harmonics of $f_{r}$ are observed in the tip region similar to case 1. However, the amplitudes at these frequencies are weaker except $3f_{r}$, which shows a slightly stronger peak (higher energy content). This could be due to the fact that the freestream velocity for case 2a was slightly higher due to the blockage caused by the grid (see figure 22 in appendix 1), therefore bringing $\lambda_{eff}$ closer to the optimum tip speed ratio (yielding maximum $C_{p}$) as can be seen from figure 15b of Biswas and Buxton [2024a]. For case 3, the harmonics such as $2f_{r}$ and $4f_{r}$ are observed, although their amplitudes are significantly diminished. $6f_{r}$ is no longer observed for case 3. Finally, for the highest $T_{i}$ case, no spectral peaks are evident other than at $f_{r}$ and $3f_{r}$ within the field of view considered here ($0.5\leq x/D\leq 5$). These frequencies can however still be present very close to the turbine blades.
Figure 4: Spectra at $x/D=4$ and $y/D=0$ for the cases (a) 1, (b) 2a, (c) 3, and (d) 4b. The transverse velocity spectra stacked in column (i) are obtained using FFT. The power spectra in the columns (ii) - (v) are obtained using Welch’s PSD estimator having window lengths (ii) 4000, (iii) 3000, (iv) 2000 and (v) 1000. The total length of the data was 5455. The windows had 50% overlap in each case.
Figures 4(i,a-d) show the transverse velocity spectra obtained using FFT at $x/D=4$ and $y/D=0$ for the same four cases of increasing turbulence intensity. The spectra obtained using FFT looked noisy and it was not directly evident if the spectra in the WMFB were discrete or continuous. Therefore, the power spectra were re-computed at the same locations using Welch’s PSD estimator to reduce the noise. The length of the overlapping windows ($w$) was varied during the process to obtain a smoother spectrum. For $w=4000$ (column ii in figure 4) and $w=3000$ (column iii in figure 4), discrete peaks are clearly observed in and around the WMFB. A smaller window size oversimplified the spectra (see columns iv - v in figure 4). Therefore, further work concerning spectral analysis around the WMFB was carried out with a window length of 3000 having a $50\%$ overlap. The no grid case showed a distinct peak in the WMFB at $St_{D}\approx 0.2$ (figure 4(a,ii-iii). Another peak was observed at $St_{D}\approx 0.4$. Although the nacelle’s shedding mode had a similar frequency [Biswas and Buxton, 2024a], it is unlikely to be sustained until $x=4D$. This frequency is therefore more likely to be a harmonic of the wake meandering mode. Note that such harmonics of the main shedding frequency are commonly observed in bluff body wakes [Biswas et al., 2022] further bolstering the link between wake meandering and a global vortex shedding mode from the turbine recently reported by Biswas and Buxton [2024a]. A discrete nature of the spectrum in the WMFB is observed even for the cases with FST, at least for the cases with low to medium $T_{i}$. For the higher $T_{i}$ cases however, the wake meandering region looked increasingly broadband likely indicating a fundamental change in the nature of the wake meandering observed. This will be discussed in more detail in the subsequent paragraphs. Interestingly, the amplitudes of the peaks in the WMFB were stronger in the presence of FST compared to the case without FST (see columns ii-iii in figure 4). This is in line with the previous studies who also reported that the amplitude of wake meandering increased in the presence of FST [Espana et al., 2011, Gambuzza and Ganapathisubramani, 2023].
Figure 5: Dominant frequency maps for the cases (a) 1, (b) 2a, (c) 3, and (d) 4b obtained based on power spectra. The power spectra are obtained using Pwelch’s PSD estimation with a window length of 3000.
4.1 Frequency maps
For a global view of the important frequencies in different parts of the wake, a dominant frequency map is created similarly to that reported in [Biswas and Buxton, 2024a]. These maps indicate the dominant frequency at a particular spatial location obtained from the power spectrum at that location. The window length used in the PSD estimation was 3000. The map of the non-turbulent case for $\lambda=6$ is shown in figure 5(a) and it looks similar to that reported in [Biswas and Buxton, 2024a] obtained through a different set of experiments. In figure 5(a), there is a small region close to the tip near the rotor where the tip vortices ($3f_{r}$) are the strongest. However, they soon interact with each other producing $f_{r}$, which remains the dominant frequency in the wake shear layer even beyond $5D$. The frequencies related to the tip vortices, especially $f_{r}$, remain dominant in a nearly axisymmetric shell type region apart from the region behind the tower [Biswas and Buxton, 2024a]. This shell type region demarcates the inner wake from the outer background fluid. In the inner wake region, multiple low frequencies dominate in large patches and these frequencies fall in the accepted frequency range of wake meandering. Among these, the largest patch has a Strouhal number $\approx 0.22$ and it initiates from near the nacelle. The low frequency zone as a whole expands radially with downstream distance.
For the cases with FST, the high frequency shell breaks down much earlier and a connection between the background turbulence and the inner wake is established in the vicinity of $x/D=2$. For the low $T_{i}$ case (figure 5(b)), although large patches of low frequencies are observed in the wake, the far wake is dominated by a single frequency $St_{D}\approx 0.29$. This frequency is also present close to the rotor in the tip region and corresponds to the frequency previously identified as $St_{D}=0.27$ using FFT in figure 3(b). The slight difference in the frequency is due to the compromised frequency resolution in the Welch’s PSD estimation. As we increase turbulence intensity, many other low frequencies appear in smaller patches, making it look more broadband and turbulent-like in the inner wake region as well as at the wake’s outer edge.
Figure 6: Power spectra in the freestream (without the turbine) and in the wake, $0.5D$ downstream of the turbine for (a) no grid case, (b) case 2a, (c) case 3 and (d) case 4b obtained using Welch’s PSD estimation with a window length of 3000.
4.2 The nature and amplitude of wake meandering in the presence of FST
To elucidate the nature of wake meandering in the presence of FST, power spectra are shown in figure 6 with and without the turbine. In each figure set, the left sub figure shows the spectrum in the freestream where the turbine would be placed. The right sub figure shows the spectrum in the turbine wake at a distance of $x/D=0.5$ from the turbine itself. The spectra correspond to the same four cases as discussed previously. The energy content in the freestream spectra is the lowest for the no grid case and it increases for the high $T_{i}$ cases as expected. For all the cases in the turbine wake, low frequency peaks can be observed in the central region (near $y=0$) that correspond to frequencies associated with the nacelle’s vortex shedding ($f_{n}$) and wake meandering ($f_{wm}$). For the no grid case, no signature of wake meandering is found in the wake near the tip region. If looked at more closely, some non-negligible energy content can be observed in the tip region for the cases 3 and 4b. Interestingly, the low $T_{i}$ case (case 2a) showed the strongest signature of wake meandering in the tip region. When considering the freestream spectrum for case 2a, a significant energy content can be observed in a similar frequency range and at a similar spatial location (see the green arrow in figure 6(b)). It appears that these frequencies in the freestream are amplified in the turbine wake. The $f_{wm}$ peak in the tip region was nearly 6 times larger than the strongest peak in the same frequency range available within the freestream for case 2a. Further, it seems that the turbine does not equally amplify all frequencies since energy is available over a broader frequency range in the freestream, while in the wake, the frequency band is narrower. A similar amplification of the freestream frequencies was also observed for the case 2b (not shown here for brevity). These observations are in line with those from recent works which suggest that the turbine has a dynamic response [Mao and Sørensen, 2018, Heisel et al., 2018, Gupta and Wan, 2019] implying some frequencies in the freestream can have a greater impact in altering the wake dynamics, or in other words, some frequencies may promote stronger wake meandering. For the medium or high $T_{i}$ cases although a significant energy content is present in the freestream in the WMFB, the amplification of wake meandering is weaker. It is worthwhile to note that FST can also indirectly induce/influence wake meandering through breaking down the tip vortices [Howard et al., 2015]. Assessing the relative importance of FST in promoting wake meandering directly or indirectly through the tip vortices is non trivial and deserves future investigations. Nevertheless, some evidence of interaction between the tip vortices and wake meandering in the presence of FST is presented in the next paragraph.
Figure 7: Spectra for case 2a at $x/D=0.5$ and (a) $y/D=0.55$ and (b) $y/D=0.53$. Subfigures (c) and (d) show the same for the case 2b.
A closer look at figure 3(b), i.e that for case 2a at $x/D=0.5$ and $y/D=0.55$, reveals a new peak at a frequency slightly higher than $3f_{r}$. To understand the origin of this peak, the spectra are reproduced on a linear scale in figure 7(a). This new frequency was found to be numerically equal to $3f_{r}+f_{wm}$, where $f_{wm}=0.27$ is the wake meandering frequency observed towards the outer edge of the wake. Similarly at $x/D=0.5$ and $y/D=0.53$, another new frequency was found which was numerically equal to $3f_{r}-f_{wm}$. A similar observation was also recorded for the case 2b as shown in figures 7(c-d). This is an excellent example of interaction between the tip vortices and wake meandering. More interestingly, the new frequencies are similar to the secondary frequencies observed in previous studies on multiscale flows involving cylinders/prisms of unequal diameters [Baj and Buxton, 2017, Cicolin et al., 2021, Biswas et al., 2022]. The secondary frequencies were numerically equal to the sum and difference of the primary shedding frequencies of the two cylinders and were shown to be produced due to the non-linear triadic interaction between the primary frequencies. Whether the same is true for the new frequencies observed here will be explored in section 6.
Figure 8: Transverse variation of PSD amplitude averaged over WMFB for different FST cases at (a) $x/D=1$, (b) $x/D=3$ and (c) $x/D=5$. Here $N_{f}$ denotes the number frequencies in the WMFB.
Figure 9: PSD amplitudes averaged over WMFB and transverse distance $y$ for different FST cases.
Next we explore how the amplitude of the spectral peaks corresponding to wake meandering vary with different FST levels. In the presence of FST, it was harder to select a single wake meandering frequency that was dominant over the entire wake region for all the cases. Therefore, the power spectrum at a spatial location was averaged over the frequencies in the WMFB to obtain an average strength of wake meandering. The nacelle’s shedding frequency at $\lambda=6$ was previously found to be around $St_{D}\approx 0.42$ [Biswas and Buxton, 2024a]. Therefore, in order to avoid unwanted contribution from the nacelle’s shedding frequency, the averaging was performed over a frequency range $0.15\leq St_{D}\leq 0.37$ (containing $N_{f}=10$ resolved frequencies) only. The average wake meandering profiles hence created are first shown at $x/D=1$ in figure 8(a) for selected cases. In the central part of the wake, i.e. in the proximity of the nacelle, the wake meandering profile for the no grid case shows the strongest peak. The amplitude thereafter decays towards the edge of the wake. This reaffirms the role of the nacelle in initiating wake meandering when the FST level is negligible. For the other cases with FST, the profiles still show a peak in the central region, although diminished, highlighting that the nacelle still has a role in inducing wake meandering in the central wake region. Unlike the no grid case however, the FST cases show significant wake meandering amplitudes towards the edge of the wake due to the amplification of frequencies from the freestream as discussed previously. Furthermore, note that towards the wake edge, the amplitude is the strongest for the low $T_{i}$ case 2a, not for case 4b which had a higher $T_{i}$ and higher $L_{v}$. This again indicates that $T_{i}$ and $L_{v}$ are not the only important parameters, rather the specific nature of the freestream spectrum is important. Further downstream (figures 8b,c), wake meandering for the no grid case gets weaker while for the other two cases it gets stronger.
To compare all the cases simultaneously, the profiles are further averaged over $y$ and the average amplitudes are shown for all the cases considered here in figure 9. Note that nearer to the turbine, for $x/D\leq 2.5$, the average wake meandering amplitudes are comparable for all the cases. Beyond $x/D=2.5$, the differences become more discernible. Wake meandering is the weakest for the no grid case. Further, the amplitude does not increase monotonically with $T_{i}$. At the farthest streamwise location, case 2a shows almost twice stronger wake meandering compared to most of the other cases.
5 Modal decomposition
The coherent motions present in a wind turbine wake have been studied using different modal decomposition techniques such as Proper Orthogonal Decomposition (POD), Dynamic Mode Decomposition (DMD) or some variant of the two [Sarmast et al., 2014, Debnath et al., 2017, De Cillis et al., 2021, Biswas and Buxton, 2024b]. These methods successfully captured the modes related to tip vortices and vortex sheddings from the tower or nacelle and wake meandering [Debnath et al., 2017, De Cillis et al., 2021, Kinjangi and Foti, 2023, Biswas and Buxton, 2024b]. For the current dataset, we use Optimal Mode Decompoisition (OMD) which is a more generalised version of the originally proposed DMD algorithm [Schmid, 2010, Wynn et al., 2013]. A detailed description of the OMD based triple decomposition technique can be found in our previous works [Baj et al., 2015, Baj and Buxton, 2017, Biswas et al., 2022, Biswas and Buxton, 2024b]. Biswas and Buxton [2024b] recently applied OMD on the wake produced from the same wind turbine model as in the present study. Various coherent modes such as the ones associated with the tip vortices, the vortex sheddings from the nacelle and the tower and wake meandering were identified and the effect of the tip speed ratio was discussed. However, the analysis was restricted to the case without any freestream turbulence. In this section, we build on the results of Biswas and Buxton [2024b] by including the effect of freestream turbulence.
5.1 Mode clustering
As discussed previously, the nature of the power spectra in the vicinity of the wake meandering frequency band (WMFB) i.e. in $0.15\leq St_{D}\leq 0.4$, changed from a more discrete nature to a more broadband nature when $T_{i}$ was increased. Due to this broadband nature of the wake meandering in the presence of FST, when OMD was performed, the FST cases showed a comparatively larger number of modes in the WMFB for which the mode shapes looked physically meaningful. An example OMD spectrum for case 2a is shown in figure 10(a). It was therefore imperative to group the spatially similar modes together and to represent them by a single representative mode. Beit-Sadi et al. [2021] recently proposed a clustering algorithm that grouped/clustered similar OMD modes together to form a smaller set of dissimilar modes. The authors used graph theory to cluster the modes in an iterative process. To assess the similarity of two modes, the authors looked at the similarity between the subspaces spanned by the real and imaginary parts of the complex OMD modes. The algorithm was later successfully applied to construct the shedding modes behind a cylinder exposed to FST [de Oliveira et al., 2025]. The procedure adopted here differs slightly from the original algorithm. The original algorithm is briefly discussed here along with the changes introduced. For each OMD mode $\phi_{l}$, Beit-Sadi et al. [2021] defined the metric $A_{l}=[\Re(\phi_{l}),\Im(\phi_{l})]$, where $\Re$ and $\Im$ denote the real and imaginary parts of $\phi_{l}$ and are stored as column vectors. To assess the similarity between modes the authors used the smallest angle between the subspaces of the modes. The smallest angle between two modes $\phi_{l}$ and $\phi_{m}$ was therefore measured by using the statistic $\theta_{lm}=\sin^{-1}\Big{(}{||A_{l}-A_{m}\Big{(}A_{l}^{T}A_{m}\Big{)}||}_{2}\Big{)}$, where ||.||||.||| | . | | denoted the largest singular value of the matrix. For simplicity, the spatial similarity was then represented by $\zeta=\cos(\theta_{lm})$.
Figure 10: $1^{st}$, $2^{nd}$ and $5^{th}$ iterations of the mode clustering algorithm for the case 2a. The chosen parameters were as follows: $r=175$, $\zeta_{spat}=0.94$, and $\zeta_{spec}=0.3$.
A spatial similarity cutoff ($\zeta_{spat}$) was defined, above which two modes could be considered similar, or dissimilar otherwise. As an example, for the spectrum shown in figure 10, clustering is performed with $\zeta_{spat}=0.94$ (note that $\zeta_{spat}=1$ means the two modes span the same subspace). This $\zeta_{spat}$ is selected based on a convergence analysis which will be discussed in the subsequent section. Additionally, a spectral similarity cutoff $\zeta_{spec}$ was imposed such that two spatially similar modes would not be clustered if their Strouhal numbers were $\zeta_{spec}$ (or more) apart. In figure 10, $\zeta_{spec}$ was set to 0.3 which is of the order of the bandwidth of the WMFB. Nonetheless, the clustering process only weakly depended on $\zeta_{spec}$ as will be shown. The spatially similar modes are then connected with an ‘edge’ (shown by the blue lines). Note from figure 10(a) that there exists a number of similar modes only in the vicinity of the WMFB. Once the similar modes are identified, the problem of clustering becomes equivalent to finding subgroups where each mode is connected with every other mode with an edge. In graph theoretic notion, such a subgroup is termed as a ‘clique’. Further, a ‘maximal-clique’ is a subgroup or clique that is not strictly contained in any other cliques (see Beit-Sadi et al. [2021] for more details).
Once the maximal cliques are obtained, each maximal clique is represented using a single representative mode. For that, each of the modes in a maximal clique are stacked in a matrix (let us denote it as $M$) as column vectors multiplied by a significance parameter $\sigma_{l}$. If the $k^{th}$ maximal clique has $n$ modes, then $M_{k}=[\sigma_{1}\phi_{1},\sigma_{2}\phi_{2}...\sigma_{n}\phi_{n}]$. Thereafter, a singular value decomposition of $M_{k}$ is performed and the first singular vector is used as the mode representing the cluster. However, it is worthwhile to note that several aspects of the method are still not strictly defined and deserve further consideration. One of them is how to assign a frequency and growth rate to a clustered mode. Further, which statistic is the best to use as $\sigma$? For example, one could project the snapshots on the OMD modes to obtain the time varying coefficients of the modes. The magnitude of the coefficient at a specific time or a norm of the coefficient could be used as $\sigma$. Here however, it was found that using such a statistic undermined the influence of the lesser damped modes within a cluster. Accordingly, a quantity related to the growth rate ($\Re(\lambda_{OMD})$) of the modes was thought to be a more reasonable measure of $\sigma$ motivating the use of the quantity $1/|\Re(\lambda_{OMD})|$. The dominant modes (growth rate closer to zero) are therefore assigned a higher significance value. This method showed faster convergence for the current data set and yielded better converged modes representing a cluster. The modes representing a cluster are also assigned a virtual growth rate and a frequency which is obtained by taking a weighted average of the frequencies and growth rates of the modes forming a cluster with $\sigma$ used as the weight. The whole process is repeated until no two modes are similar. The $1^{st}$, $2^{nd}$ and the $5^{th}$ (final) iterations for the case 2a are shown in figure 10(a-c). Note that the mode clustering algorithm significantly simplifies the spectrum in the vicinity of WMFB.
Figure 11: Variation of the number of converged modes ($N_{c}$) with $\zeta_{spat}$ for different $\zeta_{spec}$ for the cases (a) 2a, (b) 3 and (c) 4b.
The choice of the cutoff parameters $\zeta_{spat}$ and $\zeta_{spec}$ also involves some subjectivity. The effects of $\zeta_{spat}$ and $\zeta_{spec}$ on the converged number of modes ($N_{c}$) for three different cases are shown in figure 11. The rank of OMD was set to 175 (i.e. the total number of initial modes $N=175$) for all cases following our previous work [Biswas and Buxton, 2024b]. No mode reduction is achieved for $\zeta_{spat}\geq 0.96$, while for $\zeta_{spat}\leq 0.9$ the spectra are over simplified leaving too few modes. A different $\zeta_{spat}$ was selected for each case such that $N_{c}/N\approx 0.9$. For the three cases shown in figure 11, the selected cutoffs are shown by the vertical dashed line. The different symbols in figure 11 correspond to different values of $\zeta_{spec}$. Note that the mode clustering process does not depend significantly on $\zeta_{spec}$ for $\zeta_{spec}\geq 0.2$. The analysis was therefore continued with $\zeta_{spec}=0.3$ for all the cases.
Figure 12: $1^{st}$, $2^{nd}$ and final iterations of the mode clustering algorithm for the no grid case (a-c), case 3 (d-f) and case 4b (g-i).
The first, second and the final iterations for the no grid case, case 3 and case 4b are shown in figures 12. Interestingly, a significant number of similar modes are observed for the no grid case as well. However, the cluster contained a mode (with $St_{D}\approx 0.22$) that was significantly less damped compared to the other modes. This is more clearly observed in figure 12(b). Note that, the recent work by Biswas and Buxton [2024b] applied OMD on the same data set as case 1 here without mode clustering. The modes obtained with/without mode clustering were similar for the no grid case as will be discussed in the next paragraph. Mode clustering therefore simplified the spectrum in the present case, making it easier to select the physically meaningful modes which are used to form a reduced order representation of the coherent component of the flow. Further details about the mode selection process can be found in appendix 2.
Upon clustering, the selected modes are highlighted in figure 12(c) with the ‘o’ symbols. Among the tip vortex related modes, only the frequencies $f_{r}-4f_{r}$ could be captured in the OMD spectrum (figure 12(c)). The modes corresponding to $5f_{r}$ and $6f_{r}$ were therefore obtained using phase averaging, similarly to the method discussed in our previous works [Baj and Buxton, 2017, Biswas and Buxton, 2024b]. At first glance, the number of similar modes in the first iteration for the cases with FST look comparable to that for the no grid case. However, in general, equal or more iterations were required for the cases with FST to reach convergence. Furthermore, for the cases with an elevated level of FST, the clusters were not dominated by a single mode unlike case 1. For all cases, the final modes (marked by the ‘o’ symbols) were carefully selected by looking at the spatial mode shape and how the clusters evolved over iterations.
5.2 The nature of the wake meandering mode in the presence of FST
Two dominant wake meandering modes (having a frequency in the WMFB) for the four cases of increasing $T_{i}$ are shown in figure 13. For the no grid case (figure 13(a,b)), regular patterned structures of wavelength $1.5-2D$ are observed. Note that, these two modes are almost exactly the same as the modes reported by Biswas and Buxton [2024b] without mode clustering (see figures 4(a,b)), apart from a slight phase shift and a slight difference in the mode amplitude. This offers reassurance that the mode clustering algorithm adopted here does not alter the physical information embedded in the data set, rather it simplifies the analysis. For case 2a (low $T_{i}$), the dominant wake meandering modes in figures 13(c,d) still show similar regular structures, but the modes exhibit a larger transverse (along $y$) width compared to that for the no grid case. Note from figure 13(a,b) that the modes for the no grid case are not particularly energetic for $y\geq 0.5D$. The modes initiate from the vicinity of the nacelle highlighting the importance of the nacelle to providing the perturbation necessary to kick start wake meandering [Howard et al., 2015, Biswas and Buxton, 2024a]. For higher $T_{i}$, the modes show a significant energy content for $y\geq 0.5D$ as was discussed in section 4.2. The early breakdown of the tip vortices in the presence of elevated FST can provide additional perturbations in the vicinity of the outer wake. As $T_{i}$ is increased further, the modes become more irregular in nature. A higher transverse width of the modes can however still be observed, at least for the medium $T_{i}$ case (case 3).
The energy budget analysis in Biswas and Buxton [2024b] revealed that in the absence of FST, the wake meandering modes were primarily energised by the mean flow. In the presence of FST, whether the dominant energy source for the mode stays the same or not is an interesting question and it will be discussed further in the following sections.
Figure 13: Two dominant wake meandering modes for the cases (a,b) without FST, (c,d) case 2a, (e,f) case 3, (g,h) case 4b.
Figure 14: Modes associated with $3f_{r}$ (a,c,e,g) and $f_{r}$ (b,d,f,h) for the no grid case (a,b), case 2a (c,d), case 3 (e,f) and case 4b (g,h).
Figure 15: Modes associated with $3f_{r}-f_{wm}$ (a,c) and $3f_{r}+f_{wm}$ (b,d) case 2a (a,b) and case 2b (c,d).
5.3 The modes related to the tip vortices
Among the tip vortex related frequencies, only the modes associated with $3f_{r}$ and $f_{r}$ are shown in figure 14 as they were the dominant near wake structures. The modes associated with $3f_{r}$ clearly show how the tip vortices break down earlier as freestream turbulence is increased. A similar pattern is observed for the modes associated with $f_{r}$ as well as the harmonics of $f_{r}$ and $3f_{r}$ (not shown for brevity). Furthermore, the modes with FST appear to have a larger transverse spread. This is likely due to the vortex wandering or vortex jittering reported in previous studies in the presence of FST [Gambuzza and Ganapathisubramani, 2023]. Note that for cases 2a and 2b, new frequencies were observed that corresponded to the sum and difference of $3f_{r}$ and a wake meandering frequency that was dominant near the edge of the wake. These frequencies were however not present in the OMD spectrum for either of the cases. A similar observation was reported in our previous studies where the new frequencies corresponding to the sum and difference of the dominant coherent modes were not captured by OMD [Baj and Buxton, 2017, Biswas et al., 2022]. These modes were therefore obtained using phase averaging as discussed in Baj and Buxton [2017]. These are shown in figure 15 for cases 2a and 2b. The spatial orientation of the modes reveals an interesting pattern. Note from figure 6(b) that the $f_{wm}$ mode interacting with $3f_{r}$ was energetic in the freestream side ($y>0.5D$). In order to indicate the difference in the spatial location of the modes, a black line was added to figure 15 showing the loci of the points where the kinetic energy of the mode $3f_{r}$ was maximum. It can be seen that for both the cases, the high frequency mode $3f_{r}+f_{wm}$ is situated more towards the freestream side, while $3f_{r}-f_{wm}$ resided more towards the wake side. The exact same spatial orientation of the frequencies $f_{c}\pm f_{m}$ was reported by Biswas et al. [2022], Baj and Buxton [2017] for fundamentally different multiscale flows. Here, the high frequency mode resided more towards the shear layer between the two wake generating bodies, while the low frequency mode resided in the other outer side of the wake of the smaller wake generating body (having a higher shedding frequency). Such similarity between a simple two dimensional multiscale flow and a complex three dimensional multiscale flow involving rotation is truly remarkable and adds credance to the notion that these multiscale interactions are indeed universal.
Figure 16: Integrated budget terms for (a) no grid case, (b) case 2a, (c) case 3 and (d) case 4b.
6 Energy exchanges
In this section, we use the multi-scale triple-decomposed coherent kinetic energy budget equation derived by Baj and Buxton [2017] to assess the energy budgets of the individual coherent modes. Biswas and Buxton [2024b] recently used this equation to obtain the energy budget terms of each of the individual coherent modes in a wind turbine wake in the absence of FST. It was reported that the modes associated with the nacelle/tower’s vortex shedding and wake meandering were energised by the mean flow. Contrastingly, the tip vortex related modes could be energised through different pathways such as energy transfer from the mean flow or non-linear triadic interactions between various coherent modes co-existing in the wake. The role of the traidic interactions in the tip vortex stability and merging process was quantified for the first time. Yeung et al. [2024] recently proposed a Triadic Orthogonal Decomposition (TOD) which identified coherent structures that optimally account for triadic interactions and quantified the triadic energy transfers. The authors used the same PIV dataset obtained by Biswas and Buxton [2024b] and reported that the energy transfers obtained using the new method were in good agreement with Biswas and Buxton [2024b]. Although these studies have significantly improved our understanding of the origin and evolution of the various coherent structures in a wind turbine wake, the effect of freestream turbulence was not considered in these studies. This section attempts to fill that gap.
The triple-decomposed coherent kinetic energy budget equation derived by Baj and Buxton [2017] can be symbolically represented as the following:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{\partial\tilde{k}_{l}}{\partial t}=-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{C}_{l}}+\tilde{P}_{l}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{P_{l}}}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Big{(}\tilde{T}^{+}_{l}-\tilde{T}^{-}_{l}\Big{)}}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{\epsilon_{l}}}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{D_{l}}}}$ (1)
Here, the source terms on the right hand side consist of convection ($\tilde{C}_{l}$), mean flow production ($\tilde{P}_{l}$) or energy transfer from the mean flow to the $l_{th}$ coherent mode, production of stochastic turbulent kinetic energy directly from the coherent mode $l$ ($\hat{P}_{l}$), triadic energy production ($\tilde{T}^{+}_{l}-\tilde{T}^{-}_{l}$), direct dissipation from coherent mode $l$ (${\tilde{\epsilon}}_{l}$) and diffusion ($\tilde{D}_{l}$). Baj and Buxton [2017] showed that the triadic energy production term can become significant only when there exists three frequencies that linearly combine to zero or, in other words, there is a triad in the form $f_{l}\pm f_{m}\pm f_{n}=0$. The constituents of this term were defined as follows:
$\tilde{T}^{+}_{l}=-\frac{1}{2}\sum_{f_{s},f_{t}}\overline{\tilde{u}^{f_{l}}_{i}\tilde{u}^{f_{t}}_{j}\frac{\partial\tilde{u}^{f_{s}}_{i}}{\partial x_{j}}},\hskip 28.45274pt\tilde{T}^{-}_{l}=-\frac{1}{2}\sum_{f_{s},f_{t}}\overline{\tilde{u}^{f_{s}}_{i}\tilde{u}^{f_{t}}_{j}\frac{\partial\tilde{u}^{f_{l}}_{i}}{\partial x_{j}}}.$ (2)
Here, $\tilde{T}^{+}_{l}$ denotes the nonlinear energy transfer from $f_{s}$ to $f_{l}$, while $\tilde{T}^{-}_{l}$ denotes the energy lost by $f_{l}$ to $f_{s}$. In both these terms, $f_{t}$ acts as a mediator or a catalyst [Baj and Buxton, 2017, Yeung et al., 2024]. The total triadic energy gain/loss for $f_{l}$ is obtained by summing over all the frequencies ($f_{s}$, $f_{t}$) that form a triad with $f_{l}$. Note that multiple such triads can be readily identified in the tip vortex system of a wind turbine wake. Biswas and Buxton [2024b] reported these triadic interactions in the tip vortex system for a range of $\lambda$s and quantified the importance of this term on the tip vortex merging process. The other term they found to be of particular interest was the $\tilde{P}_{l}$ term since it was directly linked to wake recovery. $\tilde{P}_{l}$ is defined as $\tilde{P}_{l}=-\sum_{m}\overline{\tilde{u}^{m}_{i}\tilde{u}^{l}_{j}}\frac{\partial\overline{u}_{i}}{\partial x_{j}}$, where $m$ can be any coherent mode selected in the reduced order representation of the flow including the $l_{th}$ mode. The complete composition of each of the terms of equation 1 can be found in Baj and Buxton [2017].
Figure 17: Profiles of normalised coherent velocity correlation ($\overline{\tilde{u}^{l}\tilde{v}^{l}}$) and transverse gradient of mean streamwise velocity ($\frac{\partial\overline{u}}{\partial y}$) for the wake meandering mode extracting the highest amount of energy from the mean flow for the no grid case (a, b) and case 2a (c, d). The profiles are shown at $x/D=1$ (a, c) and $x/D=3$ (b, d).
The kinetic energy budget terms are first discussed for the four cases analysed previously. The energy budget terms are integrated over the entire domain and shown in figure 16. Since the modes related to the tip vortices show weaker energy exchanges compared to the low frequency modes (particularly for the FST cases), a separate colour bar is used for them. As a general observation, as freestream turbulence level is increased, the number of energetic low frequency modes increased. Further, the number of tip vortex related modes decreased as discussed previously. Biswas and Buxton [2024b] showed that in the presence of negligible freestream turbulence, the wake meandering mode acted like a primary mode, extracting energy from the mean flow. In the presence of FST, the mean flow production ($\tilde{P}_{l}$) term is still found to be the dominant energy source for wake meandering. The $\tilde{P}_{l}$ terms for the dominant wake meandering modes in the presence of FST were always found to be stronger than that for the no grid case which is in line with that discussed in section 4.2. Furthermore, the dominant wake meandering modes for case 2a drew the highest amount of energy from the mean flow. The $\tilde{P}_{l}$ term then became relatively weaker as $T_{i}$ was increased further, likely due to the wake meandering modes becoming less coherent (as observed in figure 13) and less efficient at extracting energy from the mean flow. This was accompanied by an enhancement in the relative contributions from the triadic interaction and diffusion term for the highest $T_{i}$ case.
The intriguing question is why is the $\tilde{P}_{l}$ term at all stronger for the wake meandering modes in the presence of FST? To answer this we need to look at the composition of the $\tilde{P}_{l}$ term. As discussed in Biswas and Buxton [2024b], utilizing the fact that the transverse gradient of streamwise mean velocity ($\frac{\partial\bar{u}}{\partial y}$) is much stronger than the other mean velocity gradients and assuming that the velocity components of the different modes are uncorrelated ($\overline{\tilde{u}^{l}\tilde{v}^{n}}\approx 0$), we can approximate the mean flow production term as $\tilde{P}_{l}\approx\overline{\tilde{u}^{l}\tilde{v}^{l}}\frac{\partial\overline{u}}{\partial y}$. This was verified to be true for all the modes and is not shown here for brevity, a more detailed discussion can be found in Biswas and Buxton [2024b]. In figure 17, the profiles of $\overline{\tilde{u}^{l}\tilde{v}^{l}}$ and $\frac{\partial\overline{u}}{\partial y}$ are shown for the wake meandering mode extracting the highest amount of energy from the mean flow for the no grid case (figure 17(a,b)) and case 2a (figure 17(c,d)), which is representative of all FST cases. These are shown at two different streamwise locations $x/D=1$ and $x/D=3$. The Strouhal numbers of the modes are $St_{D}=0.22$ (no grid case) and $St_{D}=0.37$ (case 2a). Note that, the no grid case shows the strongest gradients in the tip shear layer. However the wake meandering mode remains sheltered well within the shear layer (see figure 13(a)), inhibiting it from utilizing the stronger gradients available at the shear layer. For case 2a, the wake meandering mode extends beyond the shear layer (see figure 13(d)). As a result, although the mean velocity gradients are slightly weaker (due to enhanced turbulent mixing), the mode does a better job of exploiting this shear thereby yielding a higher rate of energy transfer from the mean flow. A similar trend was recorded for the other wake meandering modes with FST showing a higher $\tilde{P}_{l}$ term (not shown here for brevity).
Figure 18: Inter frequency triadic energy transfers for (a) no grid case, (b) case 2a and (c) case 3.
The tip vortex related modes in the presence of FST showed energy exchanges similar to the no grid case, however, the magnitudes of the terms decreased as $T_{i}$ was increased. $f_{r}$ for the no grid case was primarily energized by the mean flow production ($\tilde{P}_{l}$) term. For the FST cases, the relative contribution of the convection term for $f_{r}$ was higher than the $\tilde{P}_{l}$ term. As $T_{i}$ is increased, the location where the kinetic energy of the $f_{r}$ mode is maximum moves further upstream. Therefore a larger proportion of the $\tilde{P}_{l}$ is not captured within the PIV field of view and eventually this missing energy is ‘picked up’ by the convection term. A field of view large enough to capture the full spatial growth and decay of the mode would yield zero total convection. Note that the tip vortices ($3f_{r}$) are formed due to the difference in pressure between the pressure and suction surface of the blades [Andersen et al., 2013]. $3f_{r}$ is therefore energised upstream of the FOV and simply decays (due to different energy transfers) within the FOV under consideration. Interestingly, the $2f_{r}$ mode was found to be energized by the triadic interaction term for all the FST cases in which it could be captured. A similar observation was reported by Biswas and Buxton [2024b] when FST was negligible. $2f_{r}$ was therefore identified as a secondary mode that is produced purely due to non-linear interaction at $\lambda=6$. The magnitude of the triadic interaction term diminished drastically with $T_{i}$ indicating weaker non-linear interaction. For the highest $T_{i}$ cases, the harmonics of $f_{r}$ and $3f_{r}$ were not observed within the field of view and hence no triadic interaction could be captured. The new frequencies $3f_{r}\pm f_{wm}$ for case 2a, although weak, showed a positive contribution from the triadic interaction term, definitively establishing the fact that these are secondary modes that were produced due to the interaction between $3f_{r}$ and $f_{wm}$, similar to those observed in simplified multiscale flows [Baj and Buxton, 2017, Biswas et al., 2022].
Figure 19: (a) Streamwise variation of spanwise averaged mean flow production term for $f_{r}$ for the different FST cases. The variation of $x_{w}$ is shown against (b) $T_{i}$ and (c) $L_{v}$.
6.1 Triadic interactions
In equation 2, if we only sum over $f_{t}$, we can obtain the triadic energy transfer between $f_{l}$ and $f_{s}$. Biswas and Buxton [2024b] reported these inter-frequency triadic energy transfers in the tip vortex system (formed by the modes $f_{r}-6f_{r}$) for negligible $T_{i}$ and found the same pattern of triadic energy transfers for a range of $\lambda$s ($\lambda\geq 5.5$). Furthermore, the triadic energy transfers were mainly concentrated within the first four frequencies $f_{r}-4f_{r}$. We perform a similar analysis here for $\lambda=6$ for different levels of $T_{i}$. The inter-frequency triadic energy transfers (integrated over the entire domain) are obtained for select cases and shown in figure 18. Note that for the no grid case, the energy exchange pattern is the same as that observed for $\lambda=6$ in [Biswas and Buxton, 2024b]. Interestingly, the same pattern is observed for the low turbulence intensity cases (case 2a). However, the transfers were an order of magnitude weaker as can be seen in figure 18(b) for case 2a. It can also be seen that the new frequencies $3f_{r}\pm f_{wm}$ are energised by $3f_{r}$, i.e. the mode having the higher frequency among the two main interacting modes. This energy pathway is similar to that observed for the secondary modes in the simple multiscale flow discussed in [Biswas et al., 2022]. A similar energy exchange pattern is observed for case 3 (figure 18(c)), but the transfer magnitudes were nearly two orders of magnitude weaker in comparison to the no grid case. These results quantitatively indicate how freestream turbulence weakens the tip vortices which affects the onset of wake recovery.
6.2 Wake recovery
The tip vortices in the near wake act as a shield, restricting the exchange of momentum with the freestream [Medici, 2005, Lignarolo et al., 2015] therefore, it is essential to break them down to initiate the wake recovery process. To understand the link between the tip vortex breakdown process and wake recovery, Biswas and Buxton [2024b] studied the mean flow production term ($\tilde{P}_{l}$ in equation 1) of the tip vortex system (consisting of the frequencies $f_{r}-6f_{r}$). It was reported that the tip vortex system exchanges energy with the mean flow primarily through the turbine’s rotational frequency $f_{r}$. Nearer to the turbine, the $f_{r}$ mode extracted energy from the mean flow ($\tilde{P}_{f_{r}}>0$) and after some distance downstream, it transferred some energy back to the mean flow ($\tilde{P}_{f_{r}}<0$) therefore marking the onset location of wake recovery (henceforth denoted by $x_{w}$). This sign change in $\tilde{P}_{f_{r}}$ was shown to be due to a change in the inclination of the vortex triplet (having a frequency $f_{r}$) produced by merging [Biswas and Buxton, 2024b]. A similar conclusion was reached by Lignarolo et al. [2015] for a two bladed turbine.
In figure 19(a), we show the streamwise evolution of $\tilde{P}_{f_{r}}$, integrated along $y$, for four cases of increasing $T_{i}$. Note that, a sign change in $\tilde{P}_{f_{r}}$ is observed even with freestream turbulence. The streamwise location where it occurs moves upstream as freestream turbulence intensity is increased. This clearly shows how the wake recovers more quickly in the presence of freestream turbulence. Figures 19(b-c) show how the wake recovery onset location ($x_{w}$) changed with $T_{i}$ and $L_{v}$ respectively for all the cases. Recently, Biswas and Buxton [2024a] introduced a new length scale (termed as the convective pitch, $Lc$) which was shown to be a better length scale in the near wake compared to $D$. This length scale was defined as $Lc=\pi D/\lambda$ and could be physically interpreted as the distance traveled by a fluid parcel in the freestream in the time taken by the turbine to complete a single revolution. Here $x_{w}$ is presented in terms of $Lc{{}_{eff}}$ which is obtained based on $\lambda_{eff}$ for each case. Note that, for the cases with the highest $T_{i}$s, the $\tilde{P}_{f_{r}}$ term was negligible, at least within the field of view considered, hence, $x_{w}$ could not be measured. Nevertheless, figures 19(b-c) show some interesting trends. Apart from the fact that $x_{w}$ generally reduces with increasing FST, it appears that for low $T_{i}$, $x_{w}$ scales better with $L_{v}$. Although figure 19 lacks data for higher $L_{v}$, it is seen that for the highest $L_{v}$ for which $x_{w}$ could be calculated, $x_{w}$ clearly did not follow the decreasing trend observed for lower $L_{v}$. This indicates that there is likely a threshold length scale for which the initiation of wake recovery is the closest to the turbine. Furthermore, this threshold length scale was close to the average size of the tip vortices near the turbine. The size of the tip vortices was calculated by fitting a circle to the phase averaged vorticity field of the tip vortices ($3f_{r}$) for the no grid case (not shown here). The best fit circle had a diameter $\approx 0.07D$, close to where the trend in figure 19(c) changes. This observation is similar to that of Gambuzza and Ganapathisubramani [2023] who noted that for very large integral length scales in a non-Kolmogorov type flow, the near wake length increased back to a value close to the non-turbulent case. Note that the near wake length is loosely connected to the wake recovery location defined here as both depend on the stability of the tip vortices. Another recent study by Bourhis et al. [2025] looked at the frequency dynamics in the near wake of a model wind turbine exposed to FST. The smallest integral length scales the authors tested were comparable to the largest studied here. Unsurprisingly, they did not find any clear influence of the integral length scales on the strength of the leading tip vortex frequencies.
Figure 20: Regular grids (a) SRG38 and (b) SRG111 and fractal grids (c) SFG5 and (d) SFG7 used in this study (image reproduced from Kankanwadi [2022]).
Figure 21: (a) Mean velocity field and (b) turbulence intensity ($T_{i}$) field behind the regular grid SRG38.
7 Conclusion
The effect of freestream turbulence (FST) on the multiscale coherent dynamics in a wind turbine wake was studied using particle image velocimetry measurements. A total of 9 FST cases were studied that had varying turbulence intensity ($T_{i}$) and integral length scales ($L_{v}$). The tip vortices were found to break down closer to the turbine in the presence of FST, resulting in an earlier initiation of wake recovery [Lignarolo et al., 2015, Biswas and Buxton, 2024b]. The onset location of wake recovery ($x_{w}$) was found to depend on both $T_{i}$ and $L_{v}$. On average, it reduced monotonically with $T_{i}$. Interestingly, a stronger dependence of $x_{w}$ on $L_{v}$ was observed when $L_{v}$ was smaller than the average size of the tip vortices, likely hinting the presence of a critical $L_{v}$ beyond which the wake recovery onset location can not be pushed further upstream [Gambuzza and Ganapathisubramani, 2023, Bourhis et al., 2025].
The nature of wake meandering with and without FST was elucidated. The frequency spectra in the vicinity of the wake meandering frequency band ($0.15\leq St_{D}\leq 0.4$) changed from a more discrete nature to a more broadband nature as $T_{i}$ was increased. All the FST cases showed stronger wake meandering compared with the no FST case. However, amplitudes of wake meandering did not increase monotonically with the level of FST, highlighting the fact that the full detail of the freestream spectrum is important [Mao and Sørensen, 2018, Gupta and Wan, 2019].
The coherent modes associated with the different frequencies were studied by using Optimal Mode Decomposition (OMD). The OMD algorithm identified a larger number of modes in the vicinity of the wake meandering frequency band that were spatially (and spectrally) similar. These modes were grouped/ clustered by using the mode clustering algorithm recently introduced by Beit-Sadi et al. [2021]. The algorithm significantly simplified the OMD spectra, particularly in the vicinity of the wake meandering frequency band. As the FST level was increased, the tip vortex related modes became weaker and decayed much closer to the turbine. On the other hand, the wake meandering modes for the cases with FST were always found to be more energetic than that for the no grid case.
The energy exchanges to/from the modes were studied using the coherent kinetic energy budget equation derived by Baj and Buxton [2017]. Even with FST, the wake meandering modes were found to be primarily energised by the mean flow, although there were non-negligible contributions from the diffusion term and triadic interactions. In the presence of FST, the wake meandering modes were shown to better utilize the mean velocity gradient, yielding a higher energy transfer from the mean flow. With FST, the pattern of triadic energy exchanges in the tip vortex system remained almost identical to that reported without FST [Biswas and Buxton, 2024b]. However, the magnitudes of the energy exchanges were drastically reduced for elevated FST levels. New secondary frequencies were reported for the first time that were produced through non-linear triadic interaction between the tip vortices and wake meandering induced by FST, similarly to our previous studies on multiscale wakes [Baj and Buxton, 2017, Biswas et al., 2022].
Figure 22: Variation of mean streamwise velocity along the centreline ($y=0$, $z=0$) behind the grids (a) SRG38, (b) SRG111, (c) SFG7 and (d) SFG5. The shaded region shows the variation of mean streamwise velocity within the field of view considered.
Figure 23: Variation of turbulence intensity ($T_{i}$) at hub height with distance from the grids (a) SRG38, (b) SRG111, (c) SFG7 and (d) SFG5. The dashed vertical lines show the position of the rotor plane in different experiments. The shaded region shows the variation of $T_{i}$ within the field of view considered.
Apendix 1: Grid characterisation
A separate experimental campaign was conducted to study the effect of freestream turbulence on the wind turbine wake. To generate freestream turbulence, 4 different passive turbulence generating grids were used. These are shown in figure 20. Among these two were regular grids and two were fractal grids. The flow fields behind these grids were first characterised using PIV experiments. 2 Phantom v641 cameras were used simultaneously to give a stitched field of view length $\approx 2.45D$ ($49$cm). The FOV was $\approx 18$cm $(0.9D)$ wide and the centre of the FOV coincided with the centre of the flume (i.e. $30$cm from the bottom wall). The grid was moved upstream keeping the optical setup intact and a number of non-synchronous experiments were conducted for a particular grid to obtain time averaged statistics over a certain streamwise distance as required. As an example the mean velocity field and the turbulence intensity field behind the grid SRG38 is shown in figure 21. Note that just behind the grid, there is a significant amount of acceleration. The pressure drop across the grid reduces the water depth slightly, hence continuity yields the acceleration. Although the acceleration reduces with downstream distance, the local mean velocity is slightly greater than $U_{\infty}$ even at the farthest location measured. This increment in the freestream velocity needs to be taken into account as it can result in slight changes to the tip speed ratio of the turbine. The turbulence intensity is the highest ($\approx 14\%$) near the grid and monotonically reduces to $\approx 3.5\%$ further downstream.
The variation of the mean streamwise velocity behind all the grids is shown in figure 22. The solid black line shows the variation of the mean centreline velocity (along $y=0,z=0$). The jumps in the variation (especially for SFG7) are due to the fact that mean data from multiple non-synchronous PIV experiments were stitched together. The streamwise mean velocity also varies along the $y$ direction. Therefore, the transverse variation of the mean velocity is indicated by the shaded region. The grid is located at $x/D=0$. The dashed vertical lines show the position of the wind turbine in different experiments. Similarly, the variation of $T_{i}$ behind the grids is shown in figure 23.
Figure 24: Streamwise autocorrelation function of transverse velocity fluctuation (a) $2D$ and (b) $5D$ downstream of the grid SRG38 at the geometric centreline ($y=0,z=0$). In (b), the black curve shows the fitted bi-Gaussian curve. The strongest secondary autocorrelation peak is shown by the red ‘+’ sign. The red dashed line shows the cutoff, the data below which was discarded during fitting.
Figure 25: Variation of integral length scale ($L_{v}$) at hub height with distance from the grids (a) SRG38, (b) SRG111, (c) SFG7 and (d) SFG5. The dashed vertical lines show the position of the rotor plane in different experiments. The shaded region shows the variation of $L_{v}$ within the field of view considered.
Figure 26: The nature of the different auto-corrletaion functions (a) $2.5D$ downstream of the grid SRG38 and (b) $14D$ downstream of the grid SFG5 at the geometrical centreline ($y=0,z=0$).
Another important quantity to characterise the nature of freestream turbulence is the integral length scale ($L_{v}$). $L_{v}$ is estimated using equation 3.
[…]
Here $R_{21}^{{}^{\prime}}$ represents the streamwise autocorrelation function of transverse velocity fluctuation ($u_{2}^{\prime}$). $r\textbf{e}_{1}$ represents some displacement $r$ along the $\textbf{e}_{1}$ direction ($\textbf{e}_{1}$ is the unit vector along the streamwise direction). Finally $L_{v}$ is calculated by integrating $R_{21}^{{}^{\prime}}$ up to a certain distance $\hat{r}$ where $R_{21}^{{}^{\prime}}$ becomes zero. Measuring $\hat{r}$ is not trivial since the nature of $R_{21}^{{}^{\prime}}$ changes with streamwise distance. Figure 24 shows the variation of $R_{21}^{{}^{\prime}}$ along the the centreline ($y=0$, $z=0$), and $2D$ and $5D$ away from the grid SRG38. Closer to the grid, $R_{21}^{{}^{\prime}}$ clearly falls off to zero, while further downstream it does not. Therefore, a fitting function needs to be used. Note that the background correlation shows multiple secondary peaks, the strongest one is shown by the ‘+’ sign in figure 24(b). A cutoff twice the peak value was used (as shown by the red dashed line) and the data above the cutoff was used for fitting. A bi-Gaussian function was used for fitting (two separate Gaussian curves were fitted on both sides of the peak). As an example, the fitted curve is shown in figure figure 24(b) by the black line. Following equation 3, $L_{v}$ was calculated on both the sides and an average of the two was taken. Towards the ends of the FOV where fitting could not be performed on both sides, $L_{v}$ was estimated simply based on data from one side. Figure 25 shows the variation of integral length scale ($L_{v}$) along the geometrical centreline. Similar to figure 22, the filled contour shows transverse variation of $L_{v}$. The dashed vertical lines show the positions of the turbine in different experiments.
The integral length scales could also be obtained using other autocorrelation functions such as $R_{12}^{{}^{\prime}}$ (transverse autocorrelation function of streamwise velocity fluctuation) or $R_{11}^{{}^{\prime}}$ (streamwise autocorrelation function of streamwise velocity fluctuation). Examples of these autocorrelation functions downstream of the grids SRG38 and SFG5 are shown in figure 26. Looking at the area under the curves, one can tell that $R_{11}^{{}^{\prime}}$ yields much larger integral length scales while $R_{21}^{{}^{\prime}}$ and $R_{12}^{{}^{\prime}}$ yields integral length scales of similar size.
Figure 27: Example low frequency modes for the (a) no grid case and (b) case 4b that are not considered to be a part of the coherent component of the flow and are likely attributable to an intermittent large-scale phenomenon.
Appendix 2: Mode selection and reduced order representation of the flow
Although mode clustering simplifies the OMD spectrum, care must be taken while selecting modes from the reduced spectrum (after performing mode clustering) to represent the coherent component of the flow. The lesser damped modes generally represent a physically meaningful mode, but drawing a cutoff on the growth rate ($Re(\lambda_{OMD})$) is subjective. Looking at the nature of the spectra, the modes $Re(\lambda_{OMD})\gtrsim-0.05$, or in that vicinity, were considered. Before selecting a mode, each mode shape had to be checked manually to see if they show any physically meaningful structure. Sometimes a slightly more damped mode had to be selected since its spatial coherence was more interpretable in a physically meaningful way. Occasionally some modes were not selected despite being less damped than some other selected modes (example figure 12(i)). The spatial shape of these modes revealed very large scale smeared structures which were likely due to the intermittent passage of large scale structures having a size smaller than that indicated by the mode shapes. Example of such modes are shown in figure 27 for the no grid case and case 4b. These modes are also highlighted by a blue arrow in the spectra shown in figures 12(c) and 12(i). Furthermore, when a mode was obtained by clustering other modes, the growth rate of the new mode was simply calculated as a weighted average of the growth rates of the modes forming the cluster. This was done simply to place the new mode somewhere in the spectrum and this growth rate need not have any physical meaning. In summary, the shape of the modes involved in clustering was always checked irrespective of their growth rates.
Funding. NB gratefully acknowledges funding through the Imperial College London President’s Scholarship and the Engineering and Physical Sciences Research Council (EPSRC) through grant EP/T51780X/1. OB gratefully acknowledges funding from EPSRC through grant no. EP/V006436/1.
Declaration of interests. The authors report no conflict of interest.
For the purposes of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript (AAM) version arising.